OpenStudy (anonymous):

Please help !! describe how you would obtain the graph of the function y = f(x/3) - 4 from the graph of y =f(x).

OpenStudy (anonymous):

hey i really need help with this question please.

OpenStudy (jdoe0001):

well, you'd first need to know what f(x) is in this case we just know is a function so f(x/3) will needs to be obtained once we get that, you plug it in the 1st function

OpenStudy (anonymous):

what do you mean by f(x/3) needs to be obtained first?

OpenStudy (jdoe0001):

like for example say f(x) = 2x + 5 so f(x/3) = 2(x/3) + 5 from there then y = f(x/3) -4 now we'd know that f(x/3) = 2(x/3) + 5 so we use that y = ( 2(x/3) + 5 ) - 4

OpenStudy (anonymous):

i a sorry but I hav 0% idea on this question

OpenStudy (jdoe0001):

you evaluate the INNER function first, and expand to the OUTTER function

OpenStudy (jdoe0001):

sorta like PEMDAS you do parentheses first, then the rest

OpenStudy (anonymous):

oh wonderful thank you so much fr you help

OpenStudy (jdoe0001):

yw